The area of a rectangle is 4320 square inches. The ratio of the width to the length is 5:6. Find the length and the width. How do I do this (step by step)? Please help, thanks!
Answer by checkley77(12844)

W:L=5:6
WL=4,320
TO FIND THE L & WE NEED TO FIND A MULTIPLIER FOR THE RATIO VALUES:
5N*6N=4,320
30N^2=4,320
N62=4,320/30
N^2=144
N=SQRT144
N=12 THIS IS THE MULTIPLIER FOR THE RATIOS.
5*12=60 FOR THE WIDTH.
6*12=72 FOR THE LENGTH.
PROOF:
60*72=4,320
4,320=4,320